Upon graduating from college 40 years​ ago, Dr. Nick Riviera was already thinking of retirement. Since then he has made deposits into his retirement fund on a monthly basis in the amount of ​$130. Nick has just completed his final payment and is at last ready to retire. His retirement fund has earned 6.3​% interest compounded monthly.

a. How much has Nick accumulated in his retirement​ account?

b. In addition to all​ this, 10 years​ ago, Nick received an inheritance check for ​$20000 from his beloved uncle. He decided to deposit the entire amount into his retirement fund. What is his current balance in the​ fund?
Answer #1

Answer (a)

Monthly deposit = $130

Number of months = 40 * 12 = 480

Monthly interest rate = 6.3% /12

Amount Nick accumulated in his retirement​ account = FV(rate, nper, pmt, pv, type)

= FV (6.3%/12, 480, -130, 0, 0)

= $280,971.79

Amount Nick accumulated in his retirement​ account = $280,971.79

Answer (b)

Amount deposited = PV = $20,000

Number of months = 10 * 12 = 120

Monthly interest rate = 6.3% /12

Current balance in the​ fund = FV = PV * (1 + Periodic interest rate) ^Number of periods

= 20000 * (1 + 6.30%/12) ^120

= $37,490.38

Current balance in the​ fund = $37,490.38